Hayhayz (hayhayz):

Evaluate S5 for 300 + 150 + 75 + … and select the correct answer\below. 18.75 93.75 581.25 145.3125

Hayhayz (hayhayz):

@imqwerty

imqwerty (imqwerty):

alright what kind of series do you think it is

Hayhayz (hayhayz):

geometric series

OpenStudy (anonymous):

hi guy's

imqwerty (imqwerty):

yes correct and the formula for the sum till \(n^{th}\) term of a geometric series is this- \(\large Sum=a \left( \frac{1-r^n}{1-r} \right)\) here \(a\) is the 1st term and r is the common ratio

OpenStudy (anonymous):

gm, find, first 5 term of this series

OpenStudy (anonymous):

find common ratio, and than find term's of this series, and after this use formula as mentioned aove in @imqwerty comment

Hayhayz (hayhayz):

what would the common ratio be?

imqwerty (imqwerty):

the common ratio is the ratio of any two consecutive terms like this-> \(common ~ratio= \large \frac{2^{nd}~term}{1^{st}~term}\)

OpenStudy (anonymous):

devide first term of the series with second term, and the answer would be yours C.R.

OpenStudy (anonymous):

sorry, second by first,

OpenStudy (anonymous):

mistake, excuses

imqwerty (imqwerty):

you can take any two consecutive terms so this is also correct- \(common~ratio=\large \frac{3^{rd}~term}{2^{nd}~term}\)

Hayhayz (hayhayz):

so 0.5?

OpenStudy (anonymous):

yes right

imqwerty (imqwerty):

yes :)

Hayhayz (hayhayz):

Using the formula you put what do i plug in for the n? 5?

imqwerty (imqwerty):

yeah you need to find "\(S5\)" meaning sum till the 5th term so yeah \(n=5\) in our case :)

Hayhayz (hayhayz):

I'll tag you if I need help on more. Thanks qwerty :D

imqwerty (imqwerty):

anytime ;)